Find the volume of the solid generated by revolving the area bounded by y=4x−x2, y=x, about x=3.


1
Expert's answer
2021-03-01T06:43:13-0500

Solution.

"V=π(\int_0^3(4x-x^2)^2dx-\int_0^3x^2dx)=\newline
π(\int_0^3(16x^2-8x^3+x^4-x^2)dx)=\newline
π(\frac{16x^3}{3}-\frac{8x^4}{4}+\frac{x^5}{5}-\frac{x^3}{3})|_0^3=\newline
\frac{108}{5}π=21\frac{3}{5}π."